Need help with this accounting question, thanks!

Required information The following information applies to the questions displayed below A manufactured product has the following information for June. Actual 45, 400 lbs. $8.20 per lb Standard Direct materials Direct labor Overhead Units manufactured (5 1bs. $8 per 1b.) (3 hrs. $15 per hr.) 26,700 hrs. e $15.50 per hr. (3 hrs. e $12 per hr.) $330, 100 9,000 (1) Compute the standard cost per unit. Direct materials Direct labor Overhead Total (2) Compute the total cost variance for June. Indicate whether the cost variance is favorable or unfavorable. Total cost variance

Required information The following information applies to the questions displayed below.] A manufactured product has the following information for June. Standard Actual (5 lbs. $8 per 1b 45,400 1bs.$8.20 per lb Direct materials Direct labor Overhead Units manufactured 26,700 hrs. e s15.50 per hr. (3 hrs. $12 per hr. $330,100 9,000 Compute the direct materials price variance and the direct materials quantity variance. Indicate whether each variance is favorable or unfavorable AQ Actual Quantity SQ = Standard Quantity AP Actual Price SP Standard Price Actual Cost Standard Cost
Answer #1

Calculate standard cost per unit

Direct material (5*8) 40
Direct labor (3*15) 45
Overhead (3*12) 36
Total 121

b) Calculate total Cost variance

Standard cost = 121*9000 = $1089000

Actual Cost = 1116230

Total Cost variance = 1116230-1089000 = $27230 Unfavorable

c) Direct material variance

Actual Cost Standard Cost
AQ AP AQ*AP AQ SP AQ*SP SQ SP SQ*SP
45400 8.20 372280 45400 8 363200 45000 8 360000
9080 3200
Material price variance 9080 Unfavorable
Material quantity variance 3200 Unfavorable
Total material variance 12280 Unfavorable